# Algebroid functions on complex manifolds

### v5

@@ 4.3. Second Main Theorem for non-positive sectional curvature
Let be a nonconstant meromorphic function. Let be distinct values in Then for any there exists a subset with finite Lebesgue measure such that
@@
Similarly, we can define the defect and the simple defect of with respect to a value in
Corollary 4.14 (Defect relation).
Assume the same conditions as in Theorem If satisfies the growth condition
then
5. Nevanlinna theory of algebroid functions on Kähler manifolds
Let be a non-compact complete Kähler manifold of complex dimension with Laplace-Beltrami operator associated to its Kähler metric We adopt the same notations such as etc. defined in Section 4.
@@ 5.1. First Main Theorem
Let be an arbitrary relatively compact domain with smooth boundary in Fix a point then we have Green function for and harmonic measure on (see definition in Section 4.1).
@@
Here, are single-valued meromorphic components of namely, are locally meromorphic on outside Recall that it defines a meromorphic mapping with indeterminacy locus Suppose The characteristic function of for is defined by
@@
With the aid of Vieta’s theorem, we obtain
Using Dykin formula (see Lemma 4.1), then it yields that
From the definition, we see that the left hand side equals
@@ 5.2. Estimates of branching divisors
Recall that the discriminant of is given by
which is a -order determinant, see Section 2.1.
@@
in a local holomorphic coordinate near such that has a local coordinate If then is not a pole of One can note from an observation over that there exist terms in which contain the factor i.e., the factor appears at least times in It implies that If then is a pole of By Theorem 2.23, we have
@@ 5.3. Second Main Theorem
Let be the -leaf complex manifold of -valued algebroid function on which is a (non-compact) complex manifold of complex dimension due to Theorem 2.22. Let be the natural projection. Let be the pullback of complete Kähler metric on Note that is a complete degenerate Kähler metric on for that could be not positive-definite at points in the branching set of Using Corollary 3.8, can lift to a meromorphic function such that Note that is an analytic set of complex dimension Next, we show that Dykin formula (see Lemma 4.1) still holds under the pullback metric Set Pick a sequence of neighborhoods of such that
Equip with a complete Kähler metric Let Clearly, is a complete Kähler metric on and as Let be the Laplace-Beltrami operators of respectively. Then, weakly as Take a bounded domain with smooth boundary in Fix a point with Denote by the Green functions of for with a pole at satisfying Dirichlet boundary condition respectively, and by the harmonic measures of on with respect to respectively. It is evident that as and (weakly) uniformally on as due to
for any Borel subset of Similarly, we also have that (weakly) uniformally on as where denote the volume measures of for respectively. To unify the form, one uses to denote the volume elements of for respectively. Using Dynkin formula, it gives that
for a -class function with on outside a polar set of singularities at most. Hence, it yields immediately by letting that
This means that Dynkin formula holds under the metric Furthermore, it is not hard to check that Theorem 4.11 and Theorem 4.13 still hold for under the corresponding curvature conditions, since is an analytic set of complex dimension
@@
In the subsection, we assume that has non-negative Ricci curvature and maximal volume growth. With the pullback metric has non-negative Ricci curvature and maximal volume growth. Let be the Kähler form of Take where is a bounded domain defined in Section 4.2. Let us recall the Green function for and the harmonic measure on defined in Section 4.2. Then, the characteristic function of is defined by
@@
If is not constant, then it follows from Theorem 4.11 that for distinct values there is a subset with finite Lebesgue measure such that
@@
Let be a nonconstant -valued algebroid function on Let be distinct values in Then for any there exists a subset with finite Lebesgue measure such that
@@
Note that is a -sheeted analytic covering, and two connected components () might meet only at the branching set of and a branching point of is possibly a point in On the other hand, if is a branching point with order of we have multi-leaf function elements subordinate to equation (1), say where with and gives a branching point of with order i.e., So, it is not hard to deduce from the construction of and the definition of that
@@
Combining the above with (5.3.1) and Theorem 5.3, it yields that for any there exists a subset with finite Lebesgue measure such that
@@
We continue to treat the defect relation. Define the defect and the simple defect of with respect to respectively by
According to Theorem 5.1, we obtain The condition of maximal volume growth gives that if is nonconstant, as
Corollary 5.5 (Defect relation).
Assume the same conditions as in Theorem Then
Corollary 5.6 (Picard Theorem).
Let be a nonconstant -valued algebroid function on Then, can omit distinct values at most in
@@
In the subsection, we assume that has non-positive sectional curvature. Then, also has non-positive sectional curvature under the pullback metric Take where is the geodesic ball centered at with radius Recall the Green function for and the harmonic measure on in Section 4.3. Similarly, the characteristic function of is defined by
@@
Indeed, for
where
In order to unify the form, we write
Also, define for
Assume that We denote by the connected components of centered at respectively, here for Again, define
We see that is the positive Green function of for with a pole at satisfying Dirichlet boundary condition, and is the harmonic measure on with respect to Put
with
where is the Ricci curvature tensor on Similarly, set
where
Also, define for
If is not constant, then it follows from Theorem 4.13 that for distinct values there is a subset with finite Lebesgue measure such that
holds for and all outside
Apply the similar arguments as in the proof of Theorem 5.4, we obtain:
Theorem 5.7 (Second Main Theorem).
Let be a nonconstant -valued algebroid function on Let be distinct values in Then for any there exists a subset with finite Lebesgue measure such that
holds for all outside
Set
Corollary 5.8 (Defect relation).
Assume the same conditions as in Theorem If satisfies the growth condition then
Corollary 5.9 (Picard Theorem).
Let be a nonconstant -valued algebroid function on satisfying the growth condition then can omit distinct values at most in
6. Unicity problems of algebroid functions on Kähler manifolds
In this section, is a non-compact complete Kähler manifold of complex dimension
@@ 6.1. Propagation of algebraic dependence
Let be the Fubini-Study form on Given a -valued algebroid function on and Define
Using Dynkin formula (see Lemma 4.1), it is immediate that
On the other hand, since
then we have
@@
Let be an analytic subset of nonconstant algebroid functions on are said to be algebraically dependent on if there exists a proper indecomposable algebraic subset of such that If so, we say that are -related on
@@
We introduce the notation as follows:
If has non-negative Ricci curvature and maximal volume growth, then we denote by
@@
If has non-positive sectional curvature, then we denote by
@@
and
Let stand for the hyperplane line bundle over i.e., is the holomorphic line bundle defined by a point in It gives a holomorphic line bundle over via
@@
for where (the first Chern class of ) is a fixed positive -form.
@@
Then, it follows that
Invoking Dynkin formula (see Lemma 4.1), it is therefore
This proves the lemma. ∎
@@
with
We obtain a criteria for the propagation of algebraic dependence:
@@ 6.2. Five-value type theorem
We shall apply the propagation theory of algebraic dependence established in Section 6.1 to the unicity problems of algebroid functions on non-compact complete Kähler manifolds under two kinds of different curvature conditions.
A1 has non-negative Ricci curvature and maximal volume growth
Under this curvature condition, we show that
@@
Making use of Theorem 6.6, we have holds. Similarly, if one sets for then we obtain
which also implies holds. ∎
@@
Let be defined by (24). Set
Apply the similar arguments as in the proof of Theorem 6.7, we obtain:
@@
Let be nonconstant -valued, -valued algebroid functions on respectively. Let be distinct values in Assume that satisfy the growth condition Then
@@
The theorem can be confirmed by carrying the same arguments as in the proof of Theorem 6.7. ∎
@@
Let be nonconstant -valued algebroid functions on Assume that satisfy the growth condition If share distinct values ignoring multiplicities in then
Remark 6.11.
According to the well-known fact from the case for we see that the number of the shared values is sharp in Corollaries 6.8 and 6.10. However, Theorems 6.7 and 6.9 cannot yield the optimal number of the shared values if In doing so, we need to improve Theorems 6.7 and 6.9 so that one can derive the optimal number of the shared values.
A2 has non-negative Ricci curvature and maximal volume growth
To a general unicity problem for the value sharing ignoring multiplicities, we establish the following theorem:
@@
Let be nonconstant -valued, -valued algebroid functions on respectively. Let be distinct values in Assume that for all If satisfies
with then
@@
where It yields from Lemma 6.3 that
@@
Put together (6.2) and (6.2), it yields from Theorem 5.4 that
Therefore,
But, it contradicts with the assumption of The theorem is proved. ∎
@@
Corollary 6.13 leads to a sharp five-value type theorem:
@@
Using the similar arguments as in the proof of Theorem 6.12, we obtain:
@@
Let be nonconstant -valued, -valued algebroid functions on respectively satisfying the growth condition Let be distinct values in Assume that for all If satisfies
with then
@@
Let be nonconstant -valued, -valued algebroid functions on respectively, satisfying the growth condition Let be distinct values in Then
@@
Corollary 6.16 leads to a sharp five-value type theorem:
@@
Let be nonconstant -valued, -valued algebroid functions on respectively, satisfying the growth condition If share distinct values ignoring multiplicities in then

### v9

@@ 4.3.3. Second Main Theorem
Let be a nonconstant meromorphic function. Let be distinct values in Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside where is defined by
Proof.
Applying the similar arguments as in the proof of Theorem 4.12 and combining Theorem 4.17 with Theorem 4.19, we conclude that for any there exists a subset of finite Lebesgue measure such that
@@
In the following, we estimate When we have
This implies that the theorem holds. We consider the other case that Since as we deduce that
Thus, we conclude that
This completes the proof. ∎
5. Nevanlinna Theory of Algebroid Functions
Let be a non-compact complete Kähler manifold of with Laplace-Beltrami operator associated to the Kähler metric We utilize the same notations such as etc. defined in Section 4.
@@ 5.1. First Main Theorem
Let be an arbitrary relatively compact domain with smooth boundary in Fix a point then we have Green function for and harmonic measure on (see definitions in Section 4.1).
@@
Here, are single-valued meromorphic components of namely, are locally meromorphic on outside Recall that it defines a meromorphic mapping with indeterminacy locus We can choose such that
The characteristic function of for is defined by
@@
With the aid of Vieta’s theorem, we obtain Dynkin formula yields that
By definition, the left hand side equals
@@ 5.2. Estimates of Branching Divisors
Review that the discriminant of is given by
@@
in a local holomorphic coordinate near such that has a local coordinate If then is not a pole of So, one can derive from an observation over that there exist terms in which contain the factor i.e., the factor appears at least times in It implies that If then is a pole of By Theorem 2.22, we have
@@ 5.3. Second Main Theorem
Let be the -leaf complex manifold of a -valued algebroid function on which is a (non-compact) complex manifold of complex dimension due to Theorem 2.21. Let be the natural projection. Let be the pullback of complete Kähler metric on Note that is a complete degenerate Kähler metric on because could be not positive-definite at points in the branching set of Applying Corollary 3.8, can lift to a meromorphic function such that Note that is an analytic set of complex dimension Next, we explain that Dynkin formula (see Lemma 4.1) still holds under the pullback metric
Set Picking a sequence of neighborhoods of such that
Let be a partition of the unity subordinate to the open covering of Then, and Fix a complete Kähler metric on Let Evidently, is a complete Kähler metric on with on Since then we may assume that (otherwise, a subsequence can be chosen so that as ). Let be the Laplace-Beltrami operators of on respectively. We have weakly as Take an arbitrary bounded domain with smooth boundary in Fix a point with Denote by the Green functions of for with a pole at satisfying Dirichlet boundary condition respectively, and again by the harmonic measures of on with respect to respectively. It is not difficult to show that as and (weakly) uniformally on as due to
for any Borel subset of Similarly, we also have that (weakly) uniformally on as where denote the volume measures of for respectively. To unify the form, one uses to stand for the volume elements of for respectively. Apply Dynkin formula, it gives
for a -class function with on outside a polar set of singularities at most. Hence, it yields by letting that
This means that Dynkin formula holds under the metric Furthermore, it is not hard to check that Theorem 4.12 and Theorem 4.20 also hold for under the corresponding curvature conditions, since is an analytic set of complex dimension
@@
Assume that is non-parabolic. With the pullback metric is also non-parabolic, with non-negative Ricci curvature. Let be the Kähler form of Take where is a bounded domain defined in Section 4.2.1. Recall the Green function for and the harmonic measure on defined in Section 4.2.1. Then, the characteristic function of is defined by
@@
If is not constant, then it follows from Theorem 4.12 that for distinct values there is a subset of finite Lebesgue measure such that
@@
Let be a nonconstant -valued algebroid function on Let be distinct values in Then for any there exists a subset of finite Lebesgue measure such that
@@
Note that is a -sheeted analytic covering, and two connected components () might meet only at the branching set of and a branching point of is possibly a point in On the other hand, if is a branching point with order of we have multi-leaf function elements subordinate to equation (7), say where with and is a branching point of with order i.e., So, it is not hard to deduce from the construction of and the definition of that
@@
Combining the above with (5.3.1) and Theorem 5.3, it yields that for any there exists a subset of finite Lebesgue measure such that
@@
When we have By it is trivial to deduce Hence, we obtain:
Corollary 5.5.
Let be a nonconstant -valued algebroid function on with Let be distinct values in Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside
@@
Assume that has non-positive sectional curvature satisfying (3). Note that the curvatures of also satisfy (3) under the pullback metric In the same way, we obtain notations and by setting Apply the similar arguments as in the proof of Theorem 5.4, we can show that
Theorem 5.6 (Second Main Theorem).
Let be a nonconstant -valued algebroid function on Let be distinct values in Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside
If then and Hence, we conclude that
Corollary 5.7.
Let be a nonconstant -valued algebroid function on Let be distinct values in Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside
Let be the unit ball centered at the origin 0 in Consider equipped with the Poincaré metric
in holomorphic global coordinates where
It is known that is a non-compact complete Kähler manifold of constant sectional curvature under this metric. From (29), the associated Laplace-Beltrami operator is
Take as a reference point. The associated Poincaré distance function of from 0 on can be easily written as
Hence, we have the minimal positive global Green function of for with a pole at 0:
where denotes the standard Euclidean area of the unit sphere in By definition (see Section 2.1), we have
This gives
which is just the geodesic ball centered at 0 with radius in Hence, the boundary of is the geodesic sphere centered at 0 with radius Whence, we obtain the Green function of for with a pole at 0 satisfying Dirichlet boundary condition:
So, for with
We write
in the (geodesic) polar coordinate form, where is the Riemannian metric on induced by and
is the standard Euclidean metric on It yields from (29) that
@@
Let be the Riemannian area element of Using the second formula in (31) and volume form formula, it gives immediately
where is the standard Euclidean area element of
With these preparations, we shall give an estimate for Ricci term Let be the scalar curvature of which is a negative constant. It follows from (32) that
Substituting (30) into the expression of and calculating the integral, it is not hard to deduce that
We proceed to estimate with By
we have Hence, we obtain:
Corollary 5.8.
Let be a nonconstant -valued algebroid function on Equip with the Poincaré metric. Then
holds for all outside a subset of finite Lebesgue measure.
6. Unicity Problems of Algebroid Functions
Assume that is a non-compact complete Kähler manifold.
@@ 6.1. Propagation of Algebraic Dependence
Let be the Fubini-Study form on Given a -valued algebroid function on For define
It is not hard to deduce that Since and we obtain
@@
Let be an analytic subset of We say that nonconstant algebroid functions on are algebraically dependent on if there exists a proper indecomposable algebraic subset of such that If so, are said to be -related on
@@
We introduce the notation as follows.
When is non-parabolic with non-negative Ricci curvature and volume growth satisfying (2), denote by
@@
When has non-positive sectional curvature satisfying (3), denote by
@@
and (5) is satisfied.
Let be the hyperplane line bundle over It gives a holomorphic line bundle over
@@
where is a fixed positive -form.
@@
Thus, we have
Apply Dynkin formula, we deduce that
This completes the proof. ∎
@@
and
where
@@ 6.2. Five-Value Type Theorem
A1 has non-negative Ricci curvature
Assume that is non-parabolic with non-negative Ricci curvature and volume growth satisfying (2).
@@
By Theorem 6.6, we have holds. Similarly, when for we deduce that
which implies holds. ∎
@@
Assume that has non-positive sectional curvature satisfying (3).
@@
Let be nonconstant -valued, -valued algebroid functions on respectively. Let be distinct values in Assume that satisfy Then
@@
The argument is similar as in the proof of Theorem 6.7. ∎
@@
Let be nonconstant -valued algebroid functions on Assume that satisfy If share distinct values ignoring multiplicities in then
A2 has non-negative Ricci curvature
Assume that is non-parabolic with non-negative Ricci curvature and volume growth satisfying (2).
@@
Let be nonconstant -valued, -valued algebroid functions on respectively. Let be distinct values in Assume that for all If satisfies then
@@
It yields from Lemma 6.3 that
@@
Combining (6.2) with (6.2), it yields from Theorem 5.4 that
It is therefore
But, it contradicts with the assumption for This completes the proof. ∎
@@
Assume that has non-positive sectional curvature satisfying (3).
@@
Let be nonconstant -valued, -valued algebroid functions on satisfying respectively. Let be distinct values in Assume that for all If satisfies then
Proof.
The argument is similar as in the proof of Theorem 6.11. ∎
@@
Let be nonconstant -valued, -valued algebroid functions on satisfying respectively. Let be distinct values in Then
@@
Let be nonconstant -valued, -valued algebroid functions on satisfying respectively. If share distinct values ignoring multiplicities in then